write a quadratic equation having the given numbers as solutions -11 and -3
Answer by Cabdiqafaar(9)

Let R1= -11 and R2=-3
r1+r2=(-11)+(-3)= -14
r1+r2=(-11)*(-3)=33
x^2-(r1+r2)x+(r1*r2)=0
x^2-(-14)x+33=0
x^2+14x+33=0